Calculate the number of atoms in 1.00 g of table sugar, C12H22O11

CCalculate the number of carbon atoms in 1.00 g of table sugar


1
Expert's answer
2021-04-16T05:00:38-0400

n (C12H22O11) = m(C12H22O11)/M(C12H22O11)

nmolec= n(C12H22O11)*NA

nat= nmolec*45

nat(C)= nmolec*12


nmolec= m(C12H22O11)/M(C12H22O11) * NA = 1g / 342g/mol * 6.022*1023molec/mol = 1.76*1021 molec

nat = 1.76*1021*45= 7.92*1022

nat(C) = 1.76*1021*12= 2.11*1022


Answer: 7.92*1022 atoms, 2.11*1022 carbon atoms